OpenStudy (anonymous):

8) For the parent function y= sqrt of x, what effect does a value of k = 10 have on the graph? 9) For the parent function y=3 sqrt of x what effect does a value of a = -1 have on the graph? 10) Given h = 5 and k = 2, what is the equation of the graph if the parent function is y=3 sqrt of x? Describe the transformation. 25) For the parent function y= sqrt of x, what effect does a value of h = -3 have on the graph? 26) For the parent function y= sqrt of x, what effect does a value of h = 5 have on the graph?

OpenStudy (anonymous):

is it g(x) = f(x) + k? where f(x) is the parent function and g(x) is the resulting function after the transformation?

OpenStudy (anonymous):

for which question?

OpenStudy (anonymous):

starting at the beginning, so 8

OpenStudy (anonymous):

im going to say yes, but I dont really know for sure. so just assume so :)

OpenStudy (anonymous):

well, you should have been given some direction as to what variable represents what. typically it's like this:\[g \left( x \right)=a \cdot f \left( x-h \right)+k\]were you given something like this and direction as to what a, h and k did to the function?

OpenStudy (anonymous):

i figured out a few of them, so all I need help with now, are 9, 25, 27, and 28

OpenStudy (anonymous):

thank you so much for taking the time to help me

OpenStudy (anonymous):

no, I wasnt given anything other than what I gave you :)

OpenStudy (anonymous):

if a = -1 then all of the positive y values become negative and vise versa. what would that do to the graph?

OpenStudy (anonymous):

move it left one?

OpenStudy (anonymous):

so now all of the positive x values become negative and vise versa... what would it do to the graph?

OpenStudy (anonymous):

it would flip over

OpenStudy (anonymous):

ooops, y values, not x values.

OpenStudy (anonymous):

so that the y values are negative

OpenStudy (anonymous):